I need help solving the following non-linear systems please!!
x^2+y^2=25
y=x^2-5
Thanks!

Answer by stanbon(48581)

x^2+y^2=25
y=x^2-5
----------
rearrange 2nd equation to get:
x^2 = y+5
--------------
Substitute that into the 1st equation:
y+5 + y^2 = 25
y^2+y-20 = 0
(y+5)(y-4) = 0
y = 4 or y = -5
----------
If y = 4, x^2 = y+5 = 9; then x = +/3
If y = -5, x^2 = y+5 = 0; then x = 0
-------------------------------------
Final Solution: (0,-5), (3,4), (-3,4)
-------------------------------------
Cheers,
Stan H.